# Neural Receptive Fields, Stimulus Space Embedding and Effective Geometry of Scale-Free Networks

### v1

@@ Abstract
Understanding how neuronal dynamics couple with stimuli space and how receptive fields emerge and organize within brain networks remains a fundamental challenge in neuroscience. Several models attempted to explain these phenomena, often by adjusting the network to empirical manifestations, but struggled to achieve biological plausibility. Here, we propose a physiologically grounded model in which receptive fields and population-level attractor dynamics emerge naturally from the effective hyperbolic geometry of scale-free networks. In particular, we associate stimulus space with the boundary of a hyperbolic embedding, and study the resulting neural dynamics in both rate-based and spiking implementations. The resulting localized attractors faithfully reflect the structure of the stimulus space and capture key properties of the receptive fields without fine-tuning of local connectivity, exhibiting a direct relation between a neuron’s connectivity degree and the corresponding receptive field size. The model generalizes to stimulus spaces of arbitrary dimensionality and scale, encompassing various modalities, such as orientation and place selectivity. We also provide direct experimental evidence in support of these results, based on analyses of hippocampal place fields recorded on a linear track. Overall, our framework offers a novel organizing principle for receptive field formation and establishes a direct link between network structure, stimulus space encoding, and neural dynamics.
@@ I Introduction
Neuronal spiking dynamics is governed by receptive fields — domains within a stimulus space that elicit responses from individual cells (Fig. 1). Receptive fields, varying in size, shape, and dimensionality, are found not only in sensory systems (visual [1, 2, 3], auditory [4], olfactory [5], etc.) but also in higher-order associative networks such as in the hippocampus [7, 8], postsubiculum [9, 11], (Fig. 1a), and entorhinal cortex [12]. The entirety of receptive fields captures the structure of the stimulus space, linking it to the states of neural dynamics.
Figure 1: Neural responses are governed by the receptive fields. (a) A map of angular domains—receptive fields of head-direction cells—covering configurational space of orientations, i.e., a topological cycle. (Blue dots schematically represent the firing of individual head-direction neuron, similar to the place cell responses illustrated in panel b.) (b) Hippocampal place cells spike (red dots) when the animal appears in particular regions in the navigated environment (red circular domain). The location of each dot corresponds to the animal’s position at the time when that particular neuron fired. (c) (Bottom) Place cell receptive fields tile the navigated environment. (Top) Spiking is localized on neurons whose receptive fields intersect with the animal’s current position in space, thus reflecting the current stimulus state.
@@
The physiological basis for such behavior is believed to arise from the adaptation of the structure of neuronal connectivity to the geometry of the stimulus space manifold  [3, 6]. In particular, the phenomenon of localization arises from attractor dynamics, i.e., the nonlinear interplay between excitatory and inhibitory neuronal populations, stabilizing neuronal activity over timescales that significantly exceed the duration of an individual spike [21, 22, 23, 24]. This line of reasoning underlies most Continuous Attractor Neural Network (CANN) models, which combine local excitation with long-range or homogeneous inhibition to establish a coherent attractor dynamics [26, 28, 27, 29, 30].
@@
First, this network integrates inputs from a wide array of sensory, vestibular, motor, and associative cortices, all of which are heavily processed and interwoven within a highly intricate anatomical organization and cannot be naively metrized [41, 42, 43, 44, 45, 46]. Second, place field’s localization in the environment is largely independent from the place cells’ physical positions. Third, hippocampus exhibits a number of complex transient dynamics, e.g., new maps are generated upon exposure to each novel environment. Yet, in each map, the ongoing firing activity is localized on neurons whose place fields overlap at the animal’s current location, while other cells remain suppressed, which also points at the presence of a localized activity bump.
@@
Indeed, such a bump can be constructed by rearranging place cells into an auxiliary topographic configuration (Fig. 1c), which suggests that hippocampal activity may also be governed by CANN dynamics—provided one can identify a biologically plausible network connectivity that supports sufficient structural flexibility, define an appropriate metric on the impinging stimuli, and establish a mechanism capable of extracting their proximities in a way consistent with externally observed parameters.
@@
To address these problems, we propose an approach that circumvents the need for ad hoc adjustments of connection strengths, while enabling a principled coupling between stimulus spaces and network dynamics. The core idea is to utilize the stimulus space embedding approach and leverage implicit geometric properties of scale-free networks, an architecture that is ubiquitous across biological systems, from whole brain connectomes [47, 48, 49, 50] to neural circuits [51, 19]. A defining feature of these networks is their hierarchical connectivity—a power-law degree distribution that gives rise to hyperbolic geometric effects, without compromising structural flexibility. We use these properties to demonstrate that spiking dynamics in such networks naturally generate stable activity localization that responds adaptively to external inputs and produces complex yet faithful tilings of the stimulus space. These tilings are characterized by exponentially distributed receptive field sizes, consistent with empirical observations. Taken together, this model offers a generic and physiologically plausible framework for how diverse brain networks, including deep associative areas, may process information and produce receptive fields that effectively partition multidimensional stimulus spaces.
@@ III Results
First, when the network receives random, unstructured input, neuronal activity tends to concentrate within a dense cluster of highly connected nodes located near the center of the hyperbolic embedding. In contrast, under structured stimulation—localized within a compact angular domain— we observe that the network’s response shifts toward the periphery of the hyperbolic space, i.e., the embedding boundary, thereby mirroring the structure of the stimulus space. In other words, the activity in excitatory neurons acquires a limited angular span, i.e., formes a population activity bump localized along the angular coordinate (Fig. 3c). Multiple longitudinal tests revealed that the attractor dynamic enforces the stability of the amplitude and localization of the activity bump regardless of the stimulus size and position. Fig.  3d shows the development of activity over time for cosine-shaped stimuli of different widths and localization along the angular coordinate. In other words, we observe an emergent orientational tuning, whose mechanism, however, differs from that used in the classical ring-shaped models [32, 33, 34]. No adjustments of the synaptic connections or neighbor density alterations were used—this effect is of a purely geometro-topological nature and hence generic (Fig. 3e). The observed dynamics may correspond to a generic linearly organized firing, which may represent, e.g., a $1D$ place cell activity over a linear or circular track, or orientation-selective firing of head direction cells.
@@
The strengths of recurrent excitatory connections and peripheral inhibitions can vary within a wide but limited range without compromising structural flexibility, attractor robustness, and tractability. The system may lose attractor dynamics or become overexcited if the strength is too low or too high, typically outside physiological regimes. However, at regular connection strengths, the targeted attractor dynamics reliably emerges (see Methods, VI). More detailed investigation of the stability conditions will be provided elsewhere and is beyond the scope of this study.
@@
Next, we study the sizes of the receptive fields and their layouts by identifying the subdomains of the stimulus space that elicit reliable responses from individual neurons. Following the standard experimental methodology, we identified each field as a domain in the boundary space to which a given neuron responded with confidence, i.e., above a certain threshold. Computationally, we swept a narrow stimulus along the boundary, slowly enough to saturate the amplitude of the bump attractor in each position (see Methods, VI).
The resulting receptive fields form a dense coverage of the stimulus space (Fig. 4a), and, more importantly, exhibit exponential nesting of observed sizes(Fig. 4b). Specifically, the size of a receptive field, $s$, and the corresponding neuron’s radial coordinate, $r$, are related as
$r\propto e^{-s},$ (2)
which is consistent with the experimentally observed distribution of place field sizes [74, 75] (Fig. 4 c), and implies a direct relationship between a neuron’s receptive field size $s$ and its connectivity degree $k$,
$\log(k)\propto e^{-s}.$ (3)
The larger the receptive field of a neuron, the less its activity varies in response to small changes in stimulus.
Different neurons exhibit different scale sensitivities with respect to stimulus configuration changes. For each particular stimulus, both large and small receptive fields are engaged in driving an ensemble of active cells, descending from the upscale to the smallest ones (Fig. 4d). Consequently, the localized activity ensemble comprises neurons that span all hierarchical scales along the radial axis of the embedding. As expected, a small change in stimulus strongly affects cells with smaller receptive fields, whereas the activity of neurons with large receptive fields does not alter significantly, i.e., have different scale sensitivity (Fig. 4e). Thus, the activity of cells with large receptive fields provides a “context” for the active cells whose receptive fields are nested within them, in a sense uniting the nested cells on a larger scale [66].
@@
Spiking dynamics in hyperbolic networks also produces localized attractor dynamics. We simulated a population of $400$ excitatory and $400$ inhibitory spiking neurons, using Izhikevich’s model —a choice motivated by its numerical effectiveness and ability to produce a comprehensive scope of physiologically realistic dynamic behaviors [67]. As in the rate-based model, excitatory connections were wired into a scale-free network, while inhibitory connections were established randomly (see Methods, VI). Synaptic strengths were tuned to match the effective receptive field sizes observed in the corresponding rate-based network under identical stimulation. The external inputs were applied to the embedding $1D$ boundary , with the addition of random noise to support initial excitation.
@@
Previous analyses of these data have shown that place fields remain stable in the linear reference frame, which allows us to focus on their linear sizes [65]. As shown on Fig. 6b, place fields vary in size and their distribution is well fit by an exponential profile (Fig. 6c), i.e., indeed exhibits a hyperbolic behavior.
We further investigated the local curvature of the coverage, $\xi^{\mathrm{loc}}$, estimating from the corresponding density of receptive field sizes along the $x$-coordinate (see Methods, VI):
@@
As can be seen (Fig. 6d), pronounced peaks of $\xi^{\mathrm{loc}}$ appear at the junctions between track segments. These junctions attract the animal’s attention, consistent with natural exploratory behavior, and we thus observe a deepening of the hyperbolic representation in these regions despite the overall moderate curvature of the coverage, in agreement with previous results in large-scale environment [74]. The length of the track, around $5$ meters, being intermediate in scale between a small environment such as a laboratory cage and a large-scale open environment, provides an opportunity to obtain place fields of different sizes and to observe the emergence of low curvature hyperbolicity. Indeed, the characteristic curvature is substantially smaller ($n=253$, $\xi^{\mathrm{glb}}=-0.14$; $n=167$, $\xi^{\mathrm{glb}}=-0.13$; $n=182$, $\xi^{\mathrm{glb}}=-0.09$) than that observed in large-scale environments [74].
@@
The model also reproduces physiologically viable receptive fields in hyperbolic networks embedded to higher-dimensional hyperbolic spaces—$3D$, $4D$, etc., hyperbolic “Poincare balls” with their respective, $2D$, $3D$, etc. boundaries. Associating the latter with higher-dimensional stimuli spaces allows inducing multidimensional receptive fields occupying planar or voluminous physical spaces. We simulated such fields using equal numbers of excitatory and inhibitory neurons, scaling the overall population size to match the needs of $2D$ and $3D$ spaces (see Methods, VI).
@@
In the case of a $3D$ stimulus space, the corresponding $4D$ network embedding cannot be directly visualized. However, the model’s principles and implementation do not change: we observe the formation of distinct receptive fields for individual neurons, which collectively tile the $3D$ volume when the embedding boundary is projected onto it (Fig. 9a). Each receptive field exhibits a pronounced increase in activation frequency toward its center, as shown in the environment depicted in (Fig. 9b,c). Thus, emergence of the receptive fields from the scale-free network’s structure is a generic emergent propertie.
@@ IV Discussion
Neural dynamics across diverse brain networks is shaped by how individual neurons’ receptive fields partition the stimulus space. This organizational principle plays a central role in linking the structure of the stimulus space—whether sensory, spatial, or conceptual—to neuronal dynamics and may represent a fundamental mechanism of information encoding. The physiological mechanism underlying this phenomenon remained puzzling: how neural dynamics become functionally coupled to external stimulus spaces, and which properties of network connectivity support such mappings are questions that demand fundamental understanding. In some cases, the desired dynamics can be achieved by fine-tuning synaptic strengths according to the proximity of receptive fields and aligning network topology with that of a target manifold, as, e.g., in attractor network models of angle-selective neuronal firing. However, in deep brain networks, where inputs are highly processed and cannot be naively geometrized, the link between the synaptic architecture and the internal dynamics becomes obscured. A biologically plausible mechanism for generating receptive fields in such networks is therefore more likely to rely on a different organizing principle—one that is self-contained, more general, and independent of externally observed phenomenology.
@@
The proposed approach allows capturing the structure of stimulus space through network dynamics, via a mechanism modulated—or even induced—by the network’s connectivity. The model relies entirely on the network’s endogenous topological structure, rather than the design of synaptic efficacies based on external assessments. In particular, scale-free networks allow associating the stimulus space with their effective hyperbolic boundaries. This connection requires no assumptions about the physical arrangement of neurons—the embedding is strictly functional—and offers a natural, unifying explanation for several experimental observations, such as the appearance of low-dimensional receptive fields, their layout, the empirically observed distributions of their shapes and sizes, their network-level origin, and so forth. A curious and experimentally testable prediction of the model is that the scale of a receptive field is related to the connectivity degree of the corresponding neuron. Importantly, the scale-free property does not require locally specified connectivity; rather, it arises from a more general statistical property—the hierarchical structure of connectivity, characterized by the distribution of node degrees across the network. As long as the global degree hierarchy is maintained, the network preserves its functional properties under changes in local connectivity, synaptic plasticity, and alterations of external input. Within this framework, the stimulus space may, in principle, have arbitrary dimensionality, provided a consistent functional mapping onto the network’s effective boundary. However, the finite size of biological neural networks may impose constraints on the range of receptive field dimensionalities—a consideration that lies beyond the scope of this discussion. In this study, we maintained the functional embedding curvature close to constant [52, 53]. Natural variability of synaptic connectivity may produce higher-order effects, e.g., fluctuations of the local embedding curvatures and maximal hyperbolic embedding radii, thereby affecting the resulting neural dynamics. Experimental studies have shown that receptive fields undergo tuning and detuning across multiple timescales—an effect that may reflect a progressive deepening of the effective hyperbolic embedding in response to the intensity and duration of exploration, as well as other behavioral factors [59]. Our approach offers a framework for investigating such structural variability and its functional implications. Lastly, localized populational excitation states—bump attractors—are robust to noise at large timescales and align with the slow kinetics of external world dynamics. Such states support population-level information processing, connecting rapidly fluctuating neuronal responses with slower, mean-field ensemble dynamics. This allows a localized ensemble of excited neurons to encode the current stimulus state through a functional partition of the stimulus space with receptive fields. On the other hand, the emerging internal representations of the stimulus are also directly shaped by the receptive fields’ layout. If receptive fields are hyperbolically nested, as observed experimentally and reproduced in our model, the excited ensemble simultaneously includes neurons at multiple scales, each exhibiting different sensitivity to stimulus variations. The dynamics of neuronal responses effectively inherit the scale-free property, exhibiting sensitivity to perturbations across all spatial scales of the stimulus space. Although the role of this organization in information processing is yet to be clarified, our approach provides a means to explore it via structure-function relationships of receptive fields organization. The observation that our embedding-based framework for stimulus space, grounded in the effective geometry of scale-free networks, captures a wide range of observed phenomena strongly suggests that it may faithfully reflect the physiological mechanisms underlying the activity in various brain networks, including the hippocampus.
V Acknowledgements
We express our gratitude to Dr. Anton Chizhov for fruitful discussions.
VI Methods

### v2

@@ Abstract
Understanding how receptive fields emerge and organize within brain networks and how neural dynamics couple with stimuli space is fundamental to neuroscience. Models often rely on fine-tuning connectivity to match empirical data, which may limit biological plausibility. Here we propose a physiologically grounded alternative where receptive fields and population-level attractor dynamics arise naturally from the effective hyperbolic geometry of scale-free networks. By associating stimulus space with the boundary of a hyperbolic embedding, we simulate neural dynamics using rate-based and spiking models, revealing localized activity patterns that reflect stimulus space structure without synaptic fine-tuning. The resulting receptive fields follow experimentally observed statistics and properties, and their sizes depends on neuron’s connectivity degree. The model generalizes across stimuli dimensionalities and various modalities, such as orientation and place selectivity. Experimental analyses of hippocampal place fields recorded on a linear track support these findings. This framework offers a novel organizing principle linking network structure, stimulus space encoding, and neural dynamics, providing insights into receptive field formation across diverse brain areas.
Neural Receptive Fields; Scale-Free Networks; Hyperbolic Geometry; Stimulus Space Embedding; Neuronal Dynamics; Localized attractors
@@ I Introduction
One of the foundational questions in neuroscience concerns how neural activity becomes coupled to specific stimulus processing, how internal network dynamics relate to external variations, and how this coupling is grounded in structural organization. Across different brain regions, from primary sensory cortices to deep networks such as the hippocampus, neural responses exhibit selectivity to specific stimuli. The phenomenon is so pervasive that it is often taken as a reflection of some fundamental principle of information processing within brain networks. Yet mechanistic understanding remains largely speculative even for stimuli that are readily interpretable, as in the case of orientation encoding or spatial navigation.
Traditional models often rely on extensive fine-tuning of connectivity so that neural dynamics matches the empirically observed structure of the stimulus or responses. However, such approaches limit biological plausibility and fail to explain complex stimuli and diverse dynamical phenomena. Here we propose a different approach, introducing a distinct class of mechanisms that avoids adad hochoc constraints on local structural plasticity and is instead grounded in geometric properties of network organization at a more global level—in the statistics governing how connectivity is distributed across neurons.
Neural spiking responses across various stimuli are reflected by receptive fields — specific domains within a stimulus space that elicit activity from individual cells (Fig. 1). Receptive fields, which vary in size, shape, and dimensionality, are observed not only in sensory systems (visual [1, 2, 3], auditory [4], olfactory [5], etc.) but also in higher-order associative networks such as the hippocampus [7, 8], postsubiculum [9, 11], and entorhinal cortex (Fig. 1a). Collectively, the arrangement of receptive fields reflects the structure of the stimulus space and links it to neural dynamics.
@@
The physiological basis for such behavior is thought to arise from the adaptation of neuronal connectivity to the geometry of the stimulus space manifold [3, 6]. In particular, the phenomenon of localization is attributed to attractor dynamics—nonlinear interactions between excitatory and inhibitory neuronal populations that stabilize activity over timescales much longer than individual spikes [21, 22, 23, 24]. This reasoning forms the foundation of most Continuous Attractor Neural Network (CANN) models, which achieve coherent attractor states through a combination of local excitation and long-range or homogeneous inhibition [26, 28, 27, 29, 30].
@@
Figure 1: Neural responses are governed by the receptive fields. (a) A map of angular domains—receptive fields of head-direction cells—covering configurational space of orientations, i.e., a topological cycle. (Blue dots schematically represent the firing of individual head-direction neuron, similar to the place cell responses illustrated in panel b.) (b) Hippocampal place cells spike when the animal appears in particular regions in the navigated environment (red circular domain). Place fields tile the navigated environment and spiking is localized on neurons whose receptive fields intersect with the animal’s current position in space, thus reflecting the current stimulus state. (c) The proposed approach enables a principled coupling between stimulus spaces and network dynamics without compromising structural flexibility. The stimulus space can be associated with an effective boundary space of the network’s hyperbolic embedding due to the geometric properties of scale-free networks. Receptive-field organization inherits this hyperbolic geometry, and the resulting family of localized activity states mirrors the structure of the stimulus space.
First, this network integrates inputs from a wide array of sensory, vestibular, motor, and associative cortices, all of which are heavily processed and interwoven within a highly intricate anatomical organization and cannot be naively metrized [41, 42, 43, 44, 45, 46].
@@
Second, place field’s localization in the environment is largely independent from the place cells’ physical positions [74]. Third, hippocampus exhibits a number of complex transient dynamics, e.g., new maps are generated upon exposure to each novel environment. Yet, in each map, the ongoing firing activity is localized on neurons whose place fields overlap at the animal’s current location, while other cells remain suppressed, which also points at the presence of a localized activity bump.
@@
Indeed, such a bump can be constructed by rearranging place cells into an auxiliary topographic configuration (Fig. 1c). This suggests that hippocampal activity may also be governed by CANN dynamics, provided that one can identify a biologically plausible network connectivity with sufficient structural flexibility, define an appropriate metric on the incoming stimuli, and establish a mechanism for extracting their proximities in a manner consistent with externally observed parameters.
@@
To address these problems, we propose an approach that circumvents the need for ad hoc adjustments of local connectivity, while enabling a principled coupling between stimulus spaces and network dynamics without compromising structural flexibility. The stimulus space can be associated with an effective embedding boundary space due to the geometric properties of scale-free networks—an architecture ubiquitous across biological systems, from whole brain connectomes [47, 48, 49, 50] to neural circuits [51, 19]. A defining feature of these networks is their hierarchical connectivity, characterized by a power-law degree distribution that gives rise to hyperbolic geometric effects [52, 55]. We use these properties to demonstrate that spiking dynamics in such networks naturally generate stable activity localization that responds adaptively to external inputs and produces complex yet faithful tilings of the stimulus space. These tilings are characterized by exponentially distributed receptive field sizes, consistent with empirical observations [60].
Taken together, this model offers a framework for how scale-free network structure may support the coupling with the stimulus spaces and its partition with receptive fields, as demonstrated in our simulations and analyses of hippocampal place fields.
@@ III Results
First, when the network receives random, unstructured input, neuronal activity tends to concentrate within a dense cluster of highly connected nodes located near the center of the hyperbolic embedding. In contrast, under structured stimulation—localized within a compact angular domain— we observe that the network’s response shifts toward the periphery of the hyperbolic space, i.e., the embedding boundary, thereby mirroring the structure of the stimulus space. In other words, the activity in excitatory neurons acquires a limited angular span, i.e., formes a population activity bump localized along the angular coordinate (Fig. 3c). Multiple longitudinal tests revealed that the attractor dynamic enforces the stability of the amplitude and localization of the activity bump regardless of the stimulus size and position. Fig.  3d shows the development of activity over time for cosine-shaped stimuli of different widths and localization along the angular coordinate. In other words, we observe an emergent orientational tuning, whose mechanism, however, differs from that used in the classical ring-shaped models [32, 33, 34]. No adjustments of the synaptic connections or neighbor density alterations were used—this effect is of a purely geometro-topological nature and hence generic (Fig. 3e).
The observed dynamics may correspond to a generic linearly organized firing, which may represent, e.g., a $1D$ place cell activity over a linear or circular track, or orientation-selective firing of head direction cells.
@@
The strengths of recurrent excitatory connections and peripheral inhibitions can vary within a wide but limited range without compromising structural flexibility, attractor robustness, and tractability. The system may lose attractor dynamics or become overexcited if the strength is too low or too high, typically outside physiological regimes. However, at regular connection strengths, the targeted attractor dynamics reliably emerges (see Methods, V). More detailed investigation of the stability conditions will be provided elsewhere and is beyond the scope of this study.
@@
Next, we study the sizes of the receptive fields and their layouts by identifying the subdomains of the stimulus space that elicit reliable responses from individual neurons. Following the standard experimental methodology, we identified each field as a domain in the boundary space to which a given neuron responded with confidence, i.e., above a certain threshold. Computationally, we swept a narrow stimulus along the boundary, slowly enough to saturate the amplitude of the bump attractor in each position (see Methods, V).
The resulting receptive fields form a dense coverage of the stimulus space (Fig. 4a), and, more importantly, exhibit exponential nesting of observed sizes(Fig. 4b), which is consistent with the experimentally observed distribution of place field sizes [75, 76].
Specifically, the size of a receptive field, $S$, and the corresponding neuron’s radial coordinate, $r$ (Fig. 4 c), are related as
$S\propto e^{-\alpha r},$ (2)
and implies a direct relationship between a neuron’s receptive field size $S$ and its connectivity degree $k$,
$S\propto k^{-\beta}.$ (3)
Here $\alpha,\beta>0$ are model–dependent constants; in particular, the precise numerical value of $\beta$ is controlled both by parameters of the network model and neural dynamic model, and a detailed estimation lies beyond the scope of the present study.
The larger the receptive field of a neuron, the less its activity varies in response to small changes in stimulus. For each particular stimulus, both large and small receptive fields are engaged in driving an ensemble of active cells, descending from the upscale to the smallest ones (Fig. 4d). Consequently, the localized activity ensemble comprises neurons that span all hierarchical scales along the radial axis of the embedding. As expected, a small change in stimulus strongly affects cells with smaller receptive fields, whereas the activity of neurons with large receptive fields does not alter significantly, i.e., have different scale sensitivity (Fig. 4e). Thus, the activity of cells with large receptive fields provides a “context” for the active cells whose receptive fields are nested within them, in a sense uniting the nested cells on a larger scale [67].
@@
Spiking dynamics in hyperbolic networks also produces localized attractor dynamics. We simulated a population of $400$ excitatory and $400$ inhibitory spiking neurons, using Izhikevich’s model —a choice motivated by its numerical effectiveness and ability to produce a comprehensive scope of physiologically realistic dynamic behaviors [68]. As in the rate-based model, excitatory connections were wired into a scale-free network, while inhibitory connections were established randomly (see Methods, V). Synaptic strengths were tuned to match the effective receptive field sizes observed in the corresponding rate-based network under identical stimulation. The external inputs were applied to the embedding $1D$ boundary , with the addition of random noise to support initial excitation.
@@
Previous analyses of these data have shown that place fields remain stable in the linear reference frame, which allows us to focus on their linear sizes [66]. As shown on Fig. 6b, place fields vary in size and their distribution is well fit by an exponential profile (Fig. 6c), i.e., indeed exhibits a hyperbolic behavior. We further investigated the local curvature of the coverage, $\xi^{\mathrm{loc}}$, estimating from the corresponding density of receptive field sizes along the $x$-coordinate (see Methods, V):
@@
As can be seen (Fig. 6d), pronounced peaks of $\xi^{\mathrm{loc}}$ appear at the junctions between track segments. These junctions attract the animal’s attention, consistent with natural exploratory behavior, and we thus observe a deepening of the hyperbolic representation in these regions despite the overall moderate curvature of the coverage, in agreement with previous results in large-scale environment [75].
@@
The length of the track, around $5$ meters, being intermediate in scale between a small environment such as a laboratory cage and a large-scale open environment, provides an opportunity to obtain place fields of different sizes and to observe the emergence of low curvature hyperbolicity. Indeed, the characteristic curvature is substantially smaller ($n=253$, $\xi^{\mathrm{glb}}=-0.14$; $n=167$, $\xi^{\mathrm{glb}}=-0.13$; $n=182$, $\xi^{\mathrm{glb}}=-0.09$) than that observed in large-scale environments [75].
@@
The model also reproduces physiologically viable receptive fields in hyperbolic networks embedded in higher-dimensional hyperbolic spaces—such as $3D$- or $4D$ hyperbolic ’Poincaré balls’ with their respective boundaries. By associating these boundaries with higher-dimensional stimulus spaces, the model generates multidimensional receptive fields that occupy planar or volumetric physical spaces [76, 77, 79, 78]. We simulated such fields using equal numbers of excitatory and inhibitory neurons, scaling the overall population size to match the requirements of $2D$ and $3D$ spaces (see Methods, V).
@@
In the case of a $3D$ stimulus space, the corresponding $4D$ network embedding cannot be directly visualized. However, the model’s principles and implementation do not change: we observe the formation of distinct receptive fields for individual neurons, which collectively tile the $3D$ volume when the embedding boundary is projected onto it (Fig. 9a). Each receptive field exhibits a pronounced increase in activation frequency toward its center, as shown in the environment depicted in (Fig. 9b,c). Thus, the emergence of receptive fields from the structure of scale-free networks is a generic property of the model.
@@ IV Discussion
Neural dynamics across diverse brain networks is shaped by how individual neurons’ receptive fields partition the stimulus space [17]. This organizational principle plays a central role in linking the structure of the stimulus space—whether sensory, spatial, or conceptual—to neuronal dynamics and may represent a fundamental mechanism of information encoding. The physiological mechanism underlying this phenomenon remained puzzling: how neural dynamics become functionally coupled to external stimulus spaces, and which properties of network connectivity support such mappings are questions that demand fundamental understanding. In some cases, the desired dynamics can be achieved by fine-tuning synaptic strengths according to the proximity of receptive fields and aligning network topology with that of a target manifold, as, e.g., in attractor network models of angle-selective neuronal firing. However, in deep brain networks, where inputs are highly processed and cannot be naively geometrized, the link between the synaptic architecture and the internal dynamics becomes obscured. A biologically plausible mechanism for generating receptive fields in such networks is therefore more likely to rely on a different organizing principle—one that is self-contained, more general, and independent of externally observed phenomenology.
@@
The proposed approach allows capturing the structure of stimulus space through network dynamics, via a mechanism modulated—or even induced—by the network’s connectivity. The model relies entirely on the network’s endogenous topological structure, rather than the design of synaptic efficacies based on external assessments. In particular, scale-free networks allow associating the stimulus space with their effective hyperbolic boundaries. This connection requires no assumptions about the physical arrangement of neurons—the embedding is strictly functional—and offers a natural, unifying explanation for several experimental observations, such as the appearance of low-dimensional receptive fields, their layout, the empirically observed distributions of their shapes and sizes, their network-level origin, and so forth. A curious and experimentally testable prediction of the model is that the scale of a receptive field is related to the connectivity degree of the corresponding neuron.
Importantly, the scale-free property does not require locally specified connectivity; rather, it arises from a more general statistical property—the hierarchical structure of connectivity, characterized by the distribution of node degrees across the network. As long as the global degree hierarchy is maintained, the network preserves its functional properties under changes in local connectivity, synaptic plasticity, and alterations of external input. Within this framework, the stimulus space may, in principle, have arbitrary dimensionality, provided a consistent functional mapping onto the network’s effective boundary. However, the finite size of biological neural networks may impose constraints on the range of receptive field dimensionalities—a consideration that lies beyond the scope of this discussion.
In this study, we maintained the functional embedding curvature close to constant [52, 53]. Natural variability of synaptic connectivity may produce higher-order effects, e.g., fluctuations of the local embedding curvatures and maximal hyperbolic embedding radii, thereby affecting the resulting neural dynamics. Experimental studies have shown that receptive fields undergo tuning and detuning across multiple timescales—an effect that may reflect a progressive deepening of the effective hyperbolic embedding in response to the intensity and duration of exploration, as well as other behavioral factors [60]. Our approach offers a framework for investigating such structural variability and its functional implications.
Lastly, localized populational excitation states—bump attractors—are robust to noise at large timescales and align with the slow kinetics of external world dynamics. Such states support population-level information processing, connecting rapidly fluctuating neuronal responses with slower, mean-field ensemble dynamics. This allows a localized ensemble of excited neurons to encode the current stimulus state through a functional partition of the stimulus space with receptive fields. On the other hand, the emerging internal representations of the stimulus are also directly shaped by the receptive fields’ layout. If receptive fields are hyperbolically nested, as observed experimentally and reproduced in our model, the excited ensemble simultaneously includes neurons at multiple scales, each exhibiting different sensitivity to stimulus variations. The dynamics of neuronal responses effectively inherit the scale-free property, exhibiting sensitivity to perturbations across all spatial scales of the stimulus space. Although the role of this organization in information processing is yet to be clarified, our approach provides a means to explore it via structure-function relationships of receptive fields organization.
The observation that our embedding-based framework for stimulus space, grounded in the effective geometry of scale-free networks, captures several observed phenomena suggests that it may provide a useful model for understanding the activity in brain networks, including the hippocampus.
Acknowledgements We express our gratitude to Dr. Anton Chizhov for fruitful discussions.
V Methods